# AutoHallusion: Automatic Generation of Hallucination Benchmarks for Vision-Language Models

### v1

@@ Abstract
Large vision-language models (LVLMs) hallucinate: certain context cues in an image may trigger the language module’s overconfident and incorrect reasoning on abnormal or hypothetical objects. Though a few benchmarks have been developed to investigate LVLM hallucinations, they mainly rely on hand-crafted corner cases whose fail patterns may hardly generalize, and finetuning on them could undermine their validity. These motivate us to develop the first automatic benchmark generation approach, AutoHallusion, that harnesses a few principal strategies to create diverse hallucination examples. It probes the language modules in LVLMs for context cues and uses them to synthesize images by: (1) adding objects abnormal to the context cues; (2) for two co-occurring objects, keeping one and excluding the other; or (3) removing objects closely tied to the context cues. It then generates image-based questions whose ground-truth answers contradict the language module’s prior. A model has to overcome contextual biases and distractions to reach correct answers, while incorrect or inconsistent answers indicate hallucinations. AutoHallusion enables us to create new benchmarks at the minimum cost and thus overcomes the fragility of hand-crafted benchmarks. It also reveals common failure patterns and reasons, providing key insights to detect, avoid, or control hallucinations. Comprehensive evaluations of top-tier LVLMs, e.g., GPT-4V(ision), Gemini Pro Vision, Claude 3, and LLaVA-1.5, show a 97.7% and 98.7% success rate of hallucination induction on synthetic and real-world datasets of AutoHallusion, paving the way for a long battle against hallucinations.
@@ 1 Introduction
We evaluate SOTA LVLMs, including GPT-4V(ision), Gemini Pro Vision, Claude 3, and LLaVA-1.5, on benchmarks by AutoHallusion. It achieves success rates of 97.7% and 98.7% of inducing LVLM hallucinations on synthetic and real-world data.
@@ 2 Related Work
The recent increase in large language models (LLMs), including GPT-3 Brown et al. (2020), PaLM Chowdhery et al. (2023), and BLOOM Le Scao et al. (2023), has significantly improved natural language processing. LLaMA Touvron et al. (2023) further advanced this field, and models like Alpaca Taori et al. (2023), inspired by InstructGPT Ouyang et al. (2022) and ChatGPT, utilized human-annotated data to refine LLaMA, enhancing its interaction abilities. Additionally, Large Visual Language Models (LVLMs) such as GPT-4 Achiam et al. (2023), Flamingo Alayrac et al. (2022), Bard AI (2023), MiniGPT-4 Zhu et al. (2023), InstructBLIP Dai et al. (2024), and LLaVA Liu et al. (2024b) have developed. These models combine visual and language processing to manage both textual and visual inputs and produce textual outputs. Their architecture generally includes a visual encoder (often based on CLIP Radford et al. (2021)), a modality connection module, and an LLM. LVLMs excel in generating text descriptions from images and multi-modal learning through pre-training on image-text pairs Liu et al. (2024a) and instruction-tuning with various tasks Liu et al. (2023a); Ouyang et al. (2022); Xu et al. (2023). However, addressing hallucinations in their textual outputs remains a challenge, emphasizing the need for reliability and accuracy in real-world applications.
@@
Several benchmarks have been developed to assess hallucination in VLMs in various aspects. CHAIR Rohrbach et al. (2018) evaluates object hallucination by measuring word accuracy against ground-truth sentences and segmentation for 80 MSCOCO objects. POPE Li et al. (2023b) improves upon CHAIR for better stability and flexibility while OpenCHAIR Ben-Kish et al. (2023) extends CHAIR to open-vocabulary settings. HallusionBench Guan et al. (2024a) targets visual commonsense and reasoning with 455 visual-question control pairs. Hal-Eval Jiang et al. (2024) introduces and focuses on event hallucination while CorrelationQA Han et al. (2024) examines the impact of spurious visual inputs. Our work differs from previous benchmarks by using an auto-generated hallucination approach, synthesizing visual hallucination cases through contextual influences.
@@ 4.2.1 Abnormal Object Insertion
$q^{*}=\arg\max_{q\in\mathcal{Q}}\sum_{c\in\mathcal{C}}d(f_{\mathtt{LLM}}(\mathit{T}(c),\mathit{T}(q)),y_{q}).$ (3)
@@ 4.2.3 Correlated Object Removal
The correlated object removal strategy removes the existing object from the generated scene image $I_{s}$, while the removed object has a strong correlation with multiple contextual elements within $I_{s}$. We query such an adversary object $q^{*}$ by searching for the object with the minimum sum of distances between its language prior and the ground truth information across all contextual elements in $\mathcal{C}$:
@@ 4.3 Question Construction
In the spatial relation question, we ask about the relative positions of the target object and the scene objects. Given the bounding boxes, it’s easy to obtain the following spatial relations: Left, Right, Above, Below, Front (when the perturbed object overlaps with the scene object). Spatial relation questions are asked with multiple levels of contextual information from the image, including vanilla (no extra information), single and concatenated object-level description, and the detailed caption for the whole image, all of which are generated by the victim model.
@@ 5.1 Implementation Details
In Appendix D, we provide showcases for both data preparation methods, including the initial scene and object images and those images after manipulation, which are either generated by DALL-E-2 OpenAI (2023) or queried from the real-world image dataset.
@@ 5.3 Main Results
We have the following key observations: 1) Strategies probing inserted objects (Abnormal Object and Paired Object Insertion), achieve higher hallucination attack success rates than those probing absent objects (Correlated Object Removal strategy); 2) Questions probing the existence of objects are more effective to cause hallucinations than questions probing spatial relations; 3) GPT-4V-Turbo is the most robust to hallucination attacks among all victim LVLMs; 4) Our method achieved even higher attack success rates across all LVLMs in the real-world dataset than synthetic data. We hypothesize this comes from LVLMs lack of ability to address the complexity and diversity within the real-world data, which causes its higher vulnerability to our attack strategies when using real-world data. For more experimental results, please refer to Appendix A.
@@ Appendix B Discussion
Appendix C Question Details
@@
Appendix D More Examples
@@
Appendix E Failure Case
@@
Scene: Nighttime urban bus Detected Objects: City transit bus, City transit bus, Street light pole, Urban street tree, Bus station canopy

### v2

@@ Abstract
Large vision-language models (LVLMs) are prone to hallucinations, where certain contextual cues in an image can trigger the language module to produce overconfident and incorrect reasoning about abnormal or hypothetical objects. While some benchmarks have been developed to investigate LVLM hallucinations, they often rely on hand-crafted corner cases whose failure patterns may not generalize well. Additionally, fine-tuning on these examples could undermine their validity. To address this, we aim to scale up the number of cases through an automated approach, reducing human bias in crafting such corner cases. This motivates the development of AutoHallusion, the first automated benchmark generation approach that employs several key strategies to create a diverse range of hallucination examples. Our generated visual-question pairs pose significant challenges to LVLMs, requiring them to overcome contextual biases and distractions to arrive at correct answers. AutoHallusion enables us to create new benchmarks at the minimum cost and thus overcomes the fragility of hand-crafted benchmarks. It also reveals common failure patterns and reasons, providing key insights to detect, avoid, or control hallucinations. Comprehensive evaluations of top-tier LVLMs, e.g., GPT-4V(ision), Gemini Pro Vision, Claude 3, and LLaVA-1.5, show a 97.7% and 98.7% success rate of hallucination induction on synthetic and real-world datasets of AutoHallusion, paving the way for a long battle against hallucinations. The codebase and data can be accessed at https://github.com/wuxiyang1996/AutoHallusion.
@@ 1 Introduction
We evaluate SOTA LVLMs, including GPT-4V(ision), Gemini Pro Vision, Claude 3, and LLaVA-1.5, on benchmarks by AutoHallusion. Our method achieves success rates of 97.7% and 98.7% in inducing LVLM hallucinations on synthetic and real-world data, respectively. Based on our experiments, we curated a benchmark dataset to further evaluate model performance.
@@ 2 Related Work
The recent increase in large language models (LLMs) Liu et al. (2024c), including GPT-3 Brown et al. (2020), PaLM Chowdhery et al. (2023), and BLOOM Le Scao et al. (2023), has significantly improved natural language processing. LLaMA Touvron et al. (2023) further advanced this field, and models like Alpaca Taori et al. (2023), inspired by InstructGPT Ouyang et al. (2022) and ChatGPT, utilized human-annotated data to refine LLaMA, enhancing its interaction abilities. Additionally, Large Visual Language Models (LVLMs) such as GPT-4 Achiam et al. (2023), Flamingo Alayrac et al. (2022), Bard AI (2023), MiniGPT-4 Zhu et al. (2023), InstructBLIP Dai et al. (2024), and LLaVA Liu et al. (2024b) have developed. These models combine visual and language processing to manage both textual and visual inputs and produce textual outputs. Their architecture generally includes a visual encoder (often based on CLIP Radford et al. (2021)), a modality connection module, and an LLM. LVLMs excel in generating text descriptions from images and multi-modal learning through pre-training on image-text pairs Liu et al. (2024a) and instruction-tuning with various tasks Liu et al. (2023a); Ouyang et al. (2022); Xu et al. (2023). However, addressing hallucinations in their textual outputs remains a challenge, emphasizing the need for reliability and accuracy in real-world applications.
@@
Several benchmarks have been developed to assess hallucination in VLMs in various aspects. CHAIR Rohrbach et al. (2018) evaluates object hallucination by measuring word accuracy against ground-truth sentences and segmentation for 80 MSCOCO objects. POPE Li et al. (2023b) improves upon CHAIR for better stability and flexibility while OpenCHAIR Ben-Kish et al. (2023) extends CHAIR to open-vocabulary settings. HallusionBench Guan et al. (2024a) targets visual commonsense and reasoning with 455 visual-question control pairs. Hal-Eval Jiang et al. (2024) introduces and focuses on event hallucination while CorrelationQA Han et al. (2024) examines the impact of spurious visual inputs. Our work differs from previous benchmarks Qian et al. (2024); Wang et al. (2024) using hand-crafted examples from datasets like COCO Lin et al. (2014) or Visual Genome Krishna et al. (2017) by employing an auto-generated approach to synthesize hallucination cases through contextual influences, making our method more effective and scalable.
@@ 4.2.1 Abnormal Object Insertion
$q^{*}=\arg\max_{q\in\mathcal{Q}}\sum_{c\in\mathcal{C}}d(f_{\mathtt{LLM}}(\mathit{T}(c),\mathit{T}(q)),y_{q})$ (3)
@@ 4.2.3 Correlated Object Removal
The correlated object removal strategy removes the existing object from the generated scene image $I_{s}$, while the removed object has a strong correlation with multiple contextual elements within $I_{s}$. We use Owl-ViT Minderer et al. (2022) to extract contextual elements $\mathcal{C}$ within the image and choose the object from the detected object list as the adversary object $q^{*}$ to remove using the in-painting function from Dall-E-2 OpenAI (2023).
We choose the adversary object $q^{*}$ by searching for the object with the minimum sum of distances between its language prior and the ground truth information across all contextual elements in $\mathcal{C}$:
@@ 4.3 Question Construction
In the spatial relation question, we ask about the relative positions of the target object and the scene objects. Given the bounding boxes, it’s easy to obtain the following spatial relations: Left, Right, Above, Below, Front (when the perturbed object overlaps with the scene object). To address issues made by overlapping boxes, we remove new ones with large overlaps, retaining only the highest confidence bounding box per object. Spatial relations are determined by comparing object center positions, ensuring no overlaps and effective relation extraction. Spatial relation questions are asked with multiple levels of contextual information from the image, including vanilla (no extra information), single and concatenated object-level description, and the detailed caption for the whole image, all of which are generated by the victim model.
@@ 4.4 Hallucination Detection
We manually inspect the image-query-ground truth benchmark evaluation results by LVLMs and our inspection confirmed that, among all successful cases, 92.6% of all probing questions and their corresponding answers are correctly evaluated.
@@ 5.1 Implementation Details
In Appendix E, we provide showcases for both data preparation methods, including the initial scene and object images and those images after manipulation, which are either generated by DALL-E-2 OpenAI (2023) or queried from the real-world image dataset.
@@ 5.3 Main Results
We have the following key observations: (1) Strategies probing inserted objects (Abnormal Object and Paired Object Insertion), achieve higher hallucination attack success rates than those probing absent objects (Correlated Object Removal strategy); (2) Questions probing the existence of objects are more effective to cause hallucinations than questions probing spatial relations; (3) GPT-4V-Turbo is the most robust to hallucination attacks among all victim LVLMs; (4) Our method achieved even higher attack success rates across all LVLMs in the real-world dataset than synthetic data. We hypothesize this comes from LVLMs lack of ability to address the complexity and diversity within the real-world data, which causes its higher vulnerability to our attack strategies when using real-world data. For more experimental results, please refer to Appendix A.
@@ 5.4 Ablation Studies
Diversity (200 / 126 samples) Image Quality Effectiveness of Exi. Questions Effectiveness of Sp. Questions
Strategy Data # Scene $\uparrow$ # Obj. $\uparrow$ Origin IS $\uparrow$ Edited IS $\uparrow$ FID $\downarrow$ Overall Avg. # Q. $\downarrow$ Avg. # Q. Correctness $\downarrow$ Avg. # Q. Consistency $\downarrow$ Overall Avg. # Q. $\downarrow$ Avg. # Q. Correctness $\downarrow$ Avg. # Q. Consistency $\downarrow$
Abnormal Obj. Insertion Synthetic 152 89 4.977 $\pm$ 0.754 5.295 $\pm$ 0.988 161.56 2.162 3.243 1.622 2.318 2.134 2.537
Real-World 118 78 5.126 $\pm$ 0.715 5.143 $\pm$ 0.865 162.33 1.780 2.268 1.465 1.855 1.801 1.913
Paired Obj. Insertion Synthetic 165 70 5.936 $\pm$ 1.230 6.211 $\pm$ 1.146 152.99 2.444 3.822 1.796 2.822 2.511 3.220
Real-World 118 78 5.741 $\pm$ 0.723 5.965 $\pm$ 0.754 119.73 2.114 3.321 1.550 2.003 1.820 2.227
Correlated Obj. Removal Synthetic 193 N/A 5.455 $\pm$ 0.834 5.529 $\pm$ 0.895 220.15 3.241 2.388 4.255 1.717 1.968 1.523
Real-World 118 78 5.924 $\pm$ 0.575 5.664 $\pm$ 0.957 363.93 2.927 2.369 3.472 1.725 1.898 1.581
Table 4: Metrics to evaluate the diversity, quality, and effectiveness of the benchmark generated by our AutoHallusion. We assess the benchmark based on the following three aspects: (1) Diversity indicates the number of different scenes and objects to construct the dataset, out of 200 samples; (2) Image Quality is represented using the Inception Score Salimans et al. (2016) for both original and edited images, and Frechet Inception Distance Heusel et al. (2017) between the original and edited images; (3) Effectiveness is measured by the average number of questions constructed to induce hallucination.
Synthetic Data Real-World Data
LVLMs Overall Acc. Overall Acc. Exi. Acc. Sp. Acc. Overall Acc. Exi. Acc. Sp. Acc.
GPT-4V-Turbo Yang et al. (2023) 66.0 68.5 68.3 68.8 62.9 71.5 56.3
Gemini Pro Vision Team (2023) 51.4 53.5 59.4 43.4 48.8 70.6 31.8
Claude 3 Team (2024) 37.1 37.2 44.6 24.7 36.9 55.6 22.4
LLaVA-1.5 Liu et al. (2023c) 44.5 46.6 54.2 33.8 41.8 60.4 27.3
miniGPT4 Zhu et al. (2023) 51.0 50.2 56.4 39.7 52.1 67.7 39.9
Table 5: Leaderboard of SOTA LVLMs on the benchmark created by AutoHallusion. We highlight the top-performing model.
5.5 Benchmark Curation
Based on the data generated from our experiments and following a thorough manual inspection, we developed a benchmark for further evaluation. We generate the benchmark over all three hallucinations crafting strategies we proposed, using both scene and object images from synthetic and real-world datasets. Table 4 presents the evaluation results for all metrics. For diversity, we sampled 200 hallucination examples from the synthetic dataset and 126 from the real-world dataset. The image quality results indicate that the crafted images have quality comparable to both datasets, with FID scores showing close distribution for abnormal object insertion and paired object insertion, while object removal may cause significant variance. The effectiveness metric indicates an average of 2.5 questions to induce hallucination.
We provide a leaderboard of SOTA LVLMs on the benchmark created by AutoHallusion in Table 5, while LVLMs are evaluated on both synthetic and real-world data. Metrics include the question-answering accuracy of each LVLM, covering synthetic and real-world data as well as the existence and spatial relation questions. The evaluation results show: (1) GPT-4V-Turbo has the highest accuracy among all LVLMs evaluated; (2) All LVLMs perform better on benchmarks made by synthetic data than real-world data; (3) LVLMs are more robust to hallucinations induced by existence questions than spatial relation questions.
@@ 7 Limitation
Acknowledgments
We would like to thank Qingyuan Chen (New York University) for helpful discussions and feedback about hallucination cognition science in this paper. This material is based upon work supported by the National Science Foundation under Grant No. iis-2403436 (Boyd-Graber). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix B Discussion
Appendix C GPT4-Assisted Evaluation
We evaluated the accuracy of the LVLMs’ output by using the following prompt in GPT-4:
"Imagine you are an intelligent teacher. Thoroughly read the question, reference answer and the prediction answer to ensure a clear understanding of the information provided. Assess the correctness of the predictions. If the prediction answer does not conflict with the reference answer, please generate “correct”. If the prediction answer conflict with the reference answer, please generate “incorrect”. The output should only be “correct” or “incorrect”. The question, model generated response and ground truth is as follows: …"
Appendix D Question Details
@@
Appendix E More Examples
@@
Appendix F Failure Case
@@
Scene: Nighttime urban bus Detected Objects: City transit bus, Street light pole, Urban street tree, Bus station canopy